# Expected Free Energy-based Planning as Variational Inference

### v1

@@ Abstract
We address the problem of planning under uncertainty, where an agent must choose actions that not only achieve desired outcomes but also reduce uncertainty. Traditional methods often treat exploration and exploitation as separate objectives, lacking a unified inferential foundation. Active inference, grounded in the Free Energy Principle, offers such a foundation by minimizing Expected Free Energy (EFE)—a cost function that combines utility with epistemic drives like ambiguity resolution and novelty seeking. However, the computational burden of EFE minimization has remained a major obstacle to its scalability. In this paper, we show that EFE-based planning arises naturally from minimizing a variational free energy functional on a generative model augmented with preference and epistemic priors. This result reinforces theoretical consistency with the Free Energy Principle, by casting planning itself as variational inference. Our formulation yields optimal policies that jointly support goal achievement and information gain, while incorporating a complexity term that accounts for bounded computational resources. This unifying framework connects and extends existing methods, enabling scalable, resource-aware implementations of active inference agents.
@@ 1 Introduction
Planning under uncertainty is a fundamental challenge in both artificial intelligence and cognitive neuroscience. Agents must select actions that not only achieve desired outcomes but also reduce uncertainty about their environment. Classical approaches—rooted in reinforcement learning and optimal control—typically address this by optimizing long-term utility through value function estimation or policy learning Sutton and Barto (2018); Bertsekas (1995). However, these methods often treat reward maximization (exploitation) and uncertainty reduction (exploration) as separate objectives, using heuristics to balance them. Moreover, they struggle in high-dimensional or deep-horizon settings due to compounding errors and the curse of dimensionality.
@@ 2 The Expected Free Energy Cost Function
Policies with a lower $G(u)$ value are regarded as more favorable. Active inference processes are based on the Free Energy Principle, a sophisticated theory grounded in core physics concepts that accounts for the behavior of living systems as if $G(u)$ were their planning cost function. On a more practical level, we summarize this theory by discussing the incentives behind the three components of $G(u)$.
@@
Although $G(u)$ and $G^{\prime}(u)$ are distinct cost functions, they are centered around the same criteria. They can be regarded as specific instances of a broader policy cost function template given by
@@ 3 Related Work
Palmieri et al. (2022) introduced a comprehensive framework that unifies estimation and control through belief propagation on factor graphs, with a particular emphasis on path planning applications. Building on this perspective, van de Laar et al. (2023, 2024) extended the PAI framework by integrating epistemic value into the policy evaluation process, enabling agents to account for both expected utility and information gain during planning. Specifically, they propose modifying the Variational Free Energy (VFE) by subtracting a mutual information term when inference is performed over future (i.e., planned) segments of the factor graph. This adjustment allows reactive message passing to naturally account for both instrumental and epistemic value, yielding an interruptible and entirely local inference procedure for evaluating candidate policies.
In contrast to earlier PAI methods, the approach proposed by Van de Laar and Koudahl yields results that align with EFE minimization, but unfortunately it also introduces some conceptual and practical challenges. Conceptually, it is somewhat inelegant to alternate between different cost functions depending on the location of computation within the factor graph. This bifurcation undermines the principle of a unified objective function underlying all natural inference processes. Practically, it complicates the design and implementation of inference toolkits: developers must now account for two distinct message computations for each node—one for standard inference and another for planning—thereby significantly increasing implementation complexity and reducing modularity.
Next, we develop a PAI framework that is both consistent with the FEP and avoids the conceptual and practical limitations outlined above.
@@ 4 EFE-based Planning as Variational Inference
$\tilde{p}(u)$ $=\sigma(H[q(x|u)])$ (9a)
$\tilde{p}(x)$ $=\sigma(-H[q(y|x)])$ (9b)
$\tilde{p}(yx)$ $=\sigma(D[q(\theta|yx),q(\theta|x)])\,,$ (9c)
@@
where $G(u)$ is the expected free energy as defined in (3). In (9), $\sigma(a_{k})=\exp(a_{k})/\sum_{k^{\prime}}\exp(a_{k^{\prime}})$ is the normalized exponential (“softmax”) function, $H[\cdot]$ is the entropy functional and $D[\cdot,\cdot]$ is the Kullback-Leibler divergence (both defined in Appendix B).
@@ 5.1 Optimal Planning by Variational Inference
$\sigma(a_{k})=\frac{\exp(a_{k})}{\sum_{k^{\prime}}\exp(a_{k^{\prime}})}$ (13)
@@ 5.2 Interpretation of the Epistemic Priors
The epistemic prior $\tilde{p}(u)=\sigma(H[q(x|u)])$, introduced in (9a), imposes a bias toward selecting policies that maximize the entropy over future states $x$. This reflects an information-seeking preference, as high entropy states indicate that the agent is maintaining flexibility—keeping future states open for adaptation. Additionally, the epistemic prior $\tilde{p}(x)=\sigma(-H[q(y|x)])$ in (9b), favors policies that reduce the uncertainty over future states by selecting observations that are informative about them. Together, $\tilde{p}(u)$ and $\tilde{p}(x)$ induce a bias toward ambiguity-minimizing behavior. Similarly, the priors $\tilde{p}(u)$ and $\tilde{p}(y,x)$ jointly shape a preference for policies that maximize novelty.
@@ 5.5 Toward Scalable Synthetic Active Inference
Finally, we discuss how the results presented in this paper may pave the way for scalable and energy-efficient active inference agents. In particular, we consider how the main theorem could be extended to support message computation at the level of individual nodes within a factor graph. This discussion outlines a promising direction for future research.
@@ Appendix A Proof of the Main Theorem
$\qquad+\sum_{y\theta}q(y\theta|x)\bigg{(}\underbrace{\underbrace{-\sum_{x}q(x|u)\log q(x|u)}_{=H[q(x|u)]}-\sum_{x}q(x|u)\log\tilde{p}(u)}_{=0\text{ if }\tilde{p}(u)=\sigma(H[q(x|u)])}\bigg{)}+$
$\qquad+\sum_{x}q(x|u)\bigg{(}\underbrace{\underbrace{\sum_{y}q(y|x)\log q(y|x)}_{=-H[q(y|x)]}-\sum_{y}q(y|x)\log\tilde{p}(x)}_{=0\text{ if }\tilde{p}(x)=\sigma(-H[q(y|x)])}\bigg{)}+$
$\qquad+\sum_{yx}q(yx|u)\bigg{(}\underbrace{\underbrace{\sum_{\theta}q(\theta|yx)\log\frac{q(\theta|yx)}{q(\theta|x)}}_{D[q(\theta|yx),q(\theta|x)]}-\sum_{\theta}q(\theta|yx)\log\tilde{p}(yx)}_{=0\text{ if }\tilde{p}(yx)=\sigma(D[q(\theta|yx),q(\theta|x)])}\bigg{)}$
@@ Appendix B The Entropy and Kullback-Leibler Divergence
$\underbrace{H^{\prime}[q(x|u)]}_{\begin{subarray}{c}\text{conditional}\\ \text{entropy}\end{subarray}}\triangleq-\sum_{xu}q(xu)\log q(x|u)=E_{q(x)}\left[H[q(x|u)]\right]\,.$ (18)
@@
which is a function of $y$ and $x$. Note that (19) is not the same as, but is close to the mutual information between $y$ and $\theta$, given $x$, which is defined as

### v2

@@ Abstract
We address the problem of planning under uncertainty, where an agent must choose actions that not only achieve desired outcomes but also reduce uncertainty. Traditional methods often treat exploration and exploitation as separate objectives, lacking a unified inferential foundation. Active inference, grounded in the Free Energy Principle, provides such a foundation by minimizing Expected Free Energy (EFE), a cost function that combines utility with epistemic drives, such as ambiguity resolution and novelty seeking. However, the computational burden of EFE minimization had remained a significant obstacle to its scalability. In this paper, we show that EFE-based planning arises naturally from minimizing a variational free energy functional on a generative model augmented with preference and epistemic priors. This result reinforces theoretical consistency with the Free Energy Principle by casting planning under uncertainty itself as a form of variational inference. Our formulation yields policies that jointly support goal achievement and information gain, while incorporating a complexity term that accounts for bounded computational resources. This unifying framework connects and extends existing methods, enabling scalable, resource-aware implementations of active inference agents.
@@ 1 Introduction
Planning under uncertainty is a fundamental challenge in both artificial intelligence and cognitive neuroscience. Agents must select actions that not only achieve desired outcomes but also reduce uncertainty about their environment. Classical approaches—rooted in reinforcement learning and optimal control—typically address this by optimizing long-term utility through value function estimation or policy learning Sutton and Barto (2018); Bertsekas (2012). However, these methods often treat reward maximization (exploitation) and uncertainty reduction (exploration) as separate objectives, using heuristics to strike a balance between them. Moreover, they struggle in high-dimensional or deep-horizon settings due to compounding errors and the curse of dimensionality.
@@ 2 The Expected Free Energy Cost Function
In (3), the assumption is that $q(yx\theta u)$ is the variational posterior with respect to the joint model $p(yx\theta u)\hat{p}(x)$. In section 4, we will refine this statement. Policies with a lower $G(u)$ value are regarded as more favorable. Active inference processes are based on the Free Energy Principle, a sophisticated theory grounded in core physics concepts that accounts for the behavior of living systems as if $G(u)$ were their planning cost function. On a more practical level, we summarize this theory by discussing the incentives behind the three components of $G(u)$.
@@
Although $G(u)$ and $G^{\prime}(u)$ are distinct cost functions with different definitions of epistemic costs, they are centered around the same criteria. They can be regarded as specific instances of a broader policy cost function template given by
@@ 3 Related Work
Palmieri et al. (2022) introduced a comprehensive framework that unifies estimation and control through belief propagation on factor graphs, with a particular emphasis on path planning applications. Building on this perspective, van de Laar et al. (2024) extended the PAI framework by integrating epistemic value into the policy evaluation process, enabling agents to account for both expected utility and information gain during planning. Specifically, they propose modifying the Variational Free Energy (VFE) by subtracting a mutual information term when inference is performed over future (i.e., planned) segments of the factor graph. This adjustment allows reactive message passing to naturally account for both instrumental and epistemic value, yielding an interruptible and entirely local inference procedure for evaluating candidate policies.
In contrast to above mentioned PAI methods, the approach proposed by Van de Laar and Koudahl yields results that align with EFE minimization, but unfortunately it also introduces some conceptual and practical challenges. Conceptually, it is somewhat inelegant to alternate between different cost functions depending on the location of computation within the factor graph. This bifurcation undermines the principle of a unified objective function underlying all natural inference processes. Practically, it complicates the design and implementation of inference toolkits: developers must now account for two distinct message computations for each node—one for standard inference and another for planning—thereby significantly increasing implementation complexity and reducing modularity.
Finally, outside the FEP community and more within the reinforcement learning literature, the recent work by Lázaro-Gredilla et al. (2024) offers a compelling perspective on the relationship between planning and inference. Similar to our approach, their work highlights the role of entropy and information-seeking behavior in planning. The key distinction lies in the formulation of the inference objective: while Lázaro-Gredilla et al. (2024) demonstrate that planning corresponds to a specific weighting of entropy terms within a general variational objective, we introduce a VFE functional for a generative model that is augmented with epistemic priors, which yields EFE-based planning as a natural consequence.
In the next section, we develop a PAI framework that is not only consistent with the FEP but also addresses some of the conceptual and practical limitations of the approaches discussed above.
@@ 4 EFE-based Planning as Variational Inference
$\tilde{p}(u)$ $=\exp(H[q(x|u)])$ (9a)
$\tilde{p}(x)$ $=\exp(-H[q(y|x)])$ (9b)
$\tilde{p}(yx)$ $=\exp(D[q(\theta|yx),q(\theta|x)])\,,$ (9c)
@@
where $G(u)$ is the expected free energy as defined in (3). In (9), $H[\cdot]$ is the entropy functional and $D[\cdot,\cdot]$ is the Kullback-Leibler divergence (both defined in Appendix B).
@@
As shown in Appendix A, if one wishes to enforce normalization of the prior distributions, the epistemic prior $\tilde{p}(u)$ in (9a) can alternatively be chosen as
$\tilde{p}(u)=\sigma(H[q(x|u)])\triangleq\frac{\exp(H[q(x|u)])}{\sum_{u^{\prime}}\exp(H[q(x|u^{\prime})])},$ (11)
where $\sigma$ denotes the softmax function applied to the entropy $H[q(x|u)]$. Similar constructions apply to the priors $\tilde{p}(x)$ and $\tilde{p}(xy)$. In that case, the variational free energy $F[q]$ is shifted by a constant that does not affect the location of its minimum, and thus does not influence the outcome of the optimization.
@@ 5.1 Optimal Planning by Variational Inference
$\sigma(a)_{k}=\frac{\exp(a_{k})}{\sum_{k^{\prime}}\exp(a_{k^{\prime}})}$ (14)
@@ 5.2 Interpretation of the Epistemic Priors
The epistemic prior $\tilde{p}(u)=\exp(H[q(x|u)])$, introduced in (9a), imposes a bias toward selecting policies that maximize the entropy over future states $x$. This reflects an information-seeking preference, as high entropy states indicate that the agent is maintaining flexibility—keeping future states open for adaptation. Additionally, the epistemic prior $\tilde{p}(x)=\exp(-H[q(y|x)])$ in (9b), favors policies that reduce the uncertainty over future states by selecting observations that are informative about them. Together, $\tilde{p}(u)$ and $\tilde{p}(x)$ induce a bias toward ambiguity-minimizing behavior. Similarly, the priors $\tilde{p}(u)$ and $\tilde{p}(y,x)$ jointly shape a preference for policies that maximize novelty.
@@ 5.5 Toward Scalable Synthetic Active Inference
The results presented in this paper may also pave the way for scalable and energy-efficient active inference agents. In particular, we consider how the main theorem could be extended to support message computation at the level of individual nodes within a factor graph. This discussion outlines a promising direction for future research.
@@
A key phrase here is “react to any opportunity,” which underscores that the agent (or at a finer level of abstraction, any node in the factor graph) should only compute anything when there is an actionable opportunity to minimize free energy. If we interpret the decision to compute (and send) a variational message versus not computing (and remaining silent) as an action choice, then the EFE of these alternatives can serve as a decision criterion. A message should only be computed and sent if the EFE associated with doing so is lower than the EFE of abstaining. The EFE Theorem opens the door to evaluating the EFE of action choices via standard variational free energy minimization within an appropriately extended generative model. Thus, we foresee an energy-efficient and fully autonomous active inference process, driven solely by localized variational free energy minimization. In this framework, hand-crafted planning algorithms based on ad hoc pruning in a tree search process are replaced by an autonomously operating Bayes-optimal reactive inference process. A toolbox like RxInfer would, in principle, be capable of automating this process Bagaev et al. (2023).
5.6 Limitations
This work also raises several open questions. While, in theory, it is possible to rank policy alternatives by EFE through VFE minimization on a generative model equipped with customized epistemic priors, the current results remain conceptual: implementation details are lacking, and validation through simulations has yet to be performed. To simplify the mathematical exposition, we omitted time indices from the variables; however, incorporating time explicitly, as would be required in a full dynamical system specification, may introduce additional complexity that warrants further elaboration.
Moreover, although the epistemic priors in (9) are given in closed-form expressions, their practical implementation and online updating procedures are not straightforward. In a factor graph framework, for example, message passing through nodes representing these epistemic priors would likely require pre-computation or approximation strategies.
In summary, we view the contributions presented in this paper as a conceptual foundation for a line of research aimed at realizing PAI within active inference agents.
@@ Appendix A Proof of the Main Theorem
$\qquad+\sum_{y\theta}q(y\theta|x)\bigg{(}\underbrace{\underbrace{-\sum_{x}q(x|u)\log q(x|u)}_{=H[q(x|u)]}-\sum_{x}q(x|u)\log\tilde{p}(u)}_{=0\text{ if }\tilde{p}(u)=\exp(H[q(x|u)])}\bigg{)}$ (17f)
$\qquad+\sum_{x}q(x|u)\bigg{(}\underbrace{\underbrace{\sum_{y}q(y|x)\log q(y|x)}_{=-H[q(y|x)]}-\sum_{y}q(y|x)\log\tilde{p}(x)}_{=0\text{ if }\tilde{p}(x)=\exp(-H[q(y|x)])}\bigg{)}$
$\qquad+\sum_{yx}q(yx|u)\bigg{(}\underbrace{\underbrace{\sum_{\theta}q(\theta|yx)\log\frac{q(\theta|yx)}{q(\theta|x)}}_{D[q(\theta|yx),q(\theta|x)]}-\sum_{\theta}q(\theta|yx)\log\tilde{p}(yx)}_{=0\text{ if }\tilde{p}(yx)=\exp(D[q(\theta|yx),q(\theta|x)])}\bigg{)}$
@@
The term in brackets in 17f works out as follows:
$-\sum_{x}$ $q(x|u)\log q(x|u)-\sum_{x}q(x|u)\log\tilde{p}(u)$ (18a)
$=H[q(x|u)]-\log\tilde{p}(u)$ (18b)
=$=\begin{cases}0&\text{if }\tilde{p}(u)=\exp(H[q(x|u)])\\ \text{const}&\text{if }\tilde{p}(u)=\sigma(H[q(x|u)])\end{cases}$ (18c)
where $\text{const}=\log\left(\Sigma_{u^{\prime}}\exp(H\left[q(x|u^{\prime})\right]\right)$ and
$\sigma(H[q(x|u)])\triangleq\frac{\exp(H\left[q(x|u)\right])}{\Sigma_{u^{\prime}}\exp(H\left[q(x|u^{\prime})\right]}$ (19)
is the (normalized) softmax function. Similar derivations apply to the other epistemic priors in (9).
In the context of variational inference with variational free energy $F[q]$ as in (8), the normalization of $\tilde{p}(u)$ is inconsequential, as the additive constant does not affect the location of the minimum of $F[q]$. As long as $\tilde{p}(u)\propto\exp(H[q(x|u)])$, the results of VFE minimization will be the same.
@@ Appendix B The Entropy and Kullback-Leibler Divergence
$\underbrace{H^{\prime}[q(x|u)]}_{\begin{subarray}{c}\text{conditional}\\ \text{entropy}\end{subarray}}\triangleq-\sum_{xu}q(xu)\log q(x|u)=E_{q(u)}\left[H[q(x|u)]\right]\,.$ (21)
@@
which is a function of $y$ and $x$. Note that (22) is not the same as, but is close to the mutual information between $y$ and $\theta$, given $x$, which is a scalar value defined as